# The CLAS Collaboration

### v2

@@ III.2 Hadron identification
In Eq. (1) $\beta_{\text{n}}$ is termed the nominal value and was calculated using the particle momentum ($p_{h}$) known from the DC and the exact particle mass assumption ($m_{h}$).
Figure 4: $\beta_{h}$ versus momentum distributions for positive pion candidates. The thin black curve in the middle of the event band corresponds to the nominal $\beta_{\text{n}}$ given by Eq. (1). The red curves show the applied hadron identification cuts. Events between the red curves were treated as good pion candidates.
Experimental $\beta_{h}$ versus momentum distributions were examined for each TOF scintillator for each CLAS sector. This examination revealed that for some scintillation counters, the distributions were either shifted from their nominal positions or showed a double-band structure. To correct the timing information for such counters, a special procedure, described in Refs. Skorodumina et al. (2020); Skorodumina , was developed. In addition, a few counters were found to give unreliable signals; they were removed from consideration for both experimental data and simulation.
Figure 4 shows the $\beta_{h}$ versus momentum distribution for positive pion candidates plotted for all sectors and all reliable scintillators. The event band of pion candidates is clearly seen. The thin black curve in the middle of the event band corresponds to the nominal $\beta_{\text{n}}$ given by Eq. (1). The red curves show the applied identification cuts. Events between the red curves were selected for further analysis. Analogous identification cuts were performed for the proton and negative pion candidates.
@@ III.3 Momentum corrections
The simulation of the CLAS detector correctly propagates particles through the media and, therefore, the effect of the energy loss is already included in the efficiency and does not impact the extracted cross sections. Nevertheless, in this study, the proton momentum magnitude was corrected for the energy loss. The simulation of the CLAS detector was used to establish the correction function, which was then applied for both experimental and reconstructed Monte Carlo events Skorodumina et al. (2020); Skorodumina .
@@
Due to the undefined origin of the above effect, it cannot be simulated, and therefore, it has become conventional for CLAS data analyses to apply a special momentum correction to the experimental data. This particular study uses the electron momentum corrections that have previously been developed and tested in the analysis of the free proton part of the “e1e” dataset at the same beam energy Fedotov et al. (2017, 2018). To establish them, the approach Park et al. (2003), which was based on elastic kinematics, was used. The applied corrections include an electron momentum magnitude correction as well as an electron polar angle correction, which were developed for each CLAS sector individually.
@@ III.4.1 Fiducial cuts
The active detection solid angle of the CLAS detector was smaller than $4\pi$, in part due to the space occupied by the torus field coils. This is to say that the angles covered by the coils were not equipped with any detection system and therefore formed a “dead” area for particle detection Mecking et al. (2003). Additionally, the detection area was further limited in the polar angle from ${}^{\circ}\mathrm{}$ up to ${}^{\circ}\mathrm{}$ for electrons and up to ${}^{\circ}\mathrm{}$ for other charged particles Mecking et al. (2003).
Furthermore, as was shown in different data analyses, the edges of the active detection area also do not provide a safe region for particle reconstruction, being affected by rescattering from the coils, field distortions, and similar effects. Therefore, it has become common practice to exclude these regions from consideration by applying specific fiducial cuts. This method guarantees that events accepted in the analysis include only particles detected in “safe” areas of the detector, where the acceptance is well understood.
In this study, fiducial cuts were applied for all four final-state particles ($e^{\prime}$, $p^{\prime}$, $\pi^{+}$, and $\pi^{-}$) for both experimental events and reconstructed Monte Carlo events. The analytical shapes of the cuts are similar to those used in the analysis of the free proton part of “e1e” dataset at the same beam energy Fedotov et al. (2017, 2018), and more details can be found in Refs. Skorodumina et al. (2020); Skorodumina .
@@ III.4.2 Data quality checks
To establish data quality check cuts, the DAQ live time, as well as the yields of inclusive and elastic events normalized to the FC charge, were examined as a function of the block number. Those blocks for which these quantities demonstrated large fluctuations were excluded from the analysis. In this study, the quality check cuts are similar to those used in Refs. Fedotov et al. (2017, 2018). For more details see also Refs. Skorodumina et al. (2020); Skorodumina .
@@ III.4.3 Vertex cut
Figure 5 presents the distributions of the electron $z$-coordinate at the interaction vertex for events from full and empty target runs (black and magenta histograms, respectively). Both distributions are normalized to the corresponding charge accumulated in the Faraday cup. The value of the vertex coordinate $z$ is corrected for the effects of beam-offset [^2] The beam offset is the deviation of the beam position from the CLAS central line $(x,y)=(0,0)$ that can lead to the inaccurate determination of the vertex position. Skorodumina et al. (2020); Skorodumina . Both distributions in Fig. 5 demonstrate a well-separated peak around $z_{e^{\prime}}=2.6$ cm originating from the downstream aluminum foil. The distribution of events from the empty target runs shows two other similar peaks that correspond to the entrance and exit windows of the target cell (see also Sec. II.2).
Figure 5: Distributions of the electron $z$-coordinate at the vertex for full (black histogram) and empty (magenta histogram) target runs for CLAS sector 4. Vertical red lines show the applied cuts. Both the full and empty target distributions are normalized to the corresponding FC charge.
Empty target events were passed through the same selection procedure that was established for the liquid-deuterium data and eventually were subtracted from the latter as described in Sec. IV.5. In addition to the empty target event subtraction, a cut on the electron $z$-coordinate was applied. This cut is shown by the two vertical lines in Fig. 5: events outside these lines were excluded from the analysis.
@@ III.5.1 Reaction topologies
To identify a certain exclusive reaction, one needs to register the scattered electron and either all final hadrons or all except one. In the latter case, the four-momentum of the unregistered hadron can be deduced using energy-momentum conservation. Thus for the reaction $ep\rightarrow e^{\prime}p^{\prime}\pi^{+}\pi^{-}$ one can, in general, distinguish between four “topologies” depending on the specific combination of registered final hadrons. In this particular analysis, the following two topologies were analyzed ($X$ denotes the undetected part),
@@
  • the fully exclusive topology (all final particles registered) $ep\rightarrow e^{\prime}p^{\prime}\pi^{+}\pi^{-}X$, and
  • the $\pi^{-}$ missing topology $ep\rightarrow e^{\prime}p^{\prime}\pi^{+}X$.
@@
The statistics of the fully exclusive topology are very limited, mainly because CLAS did not cover the polar angle range $0\,^{\circ}\mathrm{}<\theta_{\textrm{lab}}<8\,^{\circ}\mathrm{}$ Mecking et al. (2003). In this experiment, the presence of this forward acceptance hole mostly impacted registration of the negative particles ($e$ and $\pi^{-}$) as their trajectories were bent by the torus magnetic field towards the beam axis. This lead to a constraint on the minimum achievable $Q^{2}$ for electrons and prevented registration of the majority of negative pions. As a consequence, the $\pi^{-}$ missing topology contains the dominant part of the statistics. The contribution of the fully exclusive topology to the total analyzed statistics varies from $\sim$5near the reaction threshold to $\sim$25at $W$ between 1.7 and 1.8 GeV.
@@
Besides the limited statistics, the fully exclusive topology also suffers from limited acceptance and therefore from a very large number of empty cells (see Sec. V.1 for more details on empty cells). These circumstances do not allow for any sensible cross section information to be obtained from this topology alone. The $\pi^{-}$ missing topology has a tolerable number of empty cells and large statistics, and therefore serves the purpose of the cross section extraction best.
@@
For both analyzed topologies, a small admixture of the three-pion background is present. This is similar to the double-pion production off free protons, where the main background channel is $ep\rightarrow e^{\prime}p^{\prime}\pi^{+}\pi^{-}\pi^{0}$. The free-proton analysis Fedotov et al. (2017, 2018) that was carried out for the same beam energy $E_{beam}=2.039$ GeV demonstrated that although the admixture of the events from this background channel becomes discernible at $W\gtrsim 1.6$ GeV, it remains negligible and well separated from the double-pion events via the exclusivity cuts. For the double-pion production off protons in deuterium, one more background channel can be distinguished, which is $en(p)\rightarrow e^{\prime}p^{\prime}(p^{\prime})\pi^{+}\pi^{-}\pi^{-}$, but background events from this channel follow the same kinematic pattern as events from the aforementioned $ep\rightarrow e^{\prime}p^{\prime}\pi^{+}\pi^{-}\pi^{0}$ reaction.
@@
In general, two more topologies can be distinguished, i.e. the proton missing topology and the $\pi^{+}$ missing topology. Both require registration of the $\pi^{-}$ in the final state and as a result suffer from similar issues of suppressed statistics and limited acceptance as the fully exclusive topology. These two topologies are typically ignored in double-pion analyses Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017). Nevertheless, as demonstrated in a previous free-proton analysis Fedotov et al. (2017, 2018), all four reaction topologies can be used in combination, which allows for an increase in the statistics and a reduction in the number of empty cells.
@@
However, if the pion pair was produced off the proton bound in deuterium, these two additional topologies turn out to be contaminated with events from other reactions. Specifically, in the proton missing topology, the missing particle reconstruction fails to determine whether the pion pair was produced off the proton or off the neutron because their masses are almost identical. A similar situation occurs for the $\pi^{+}$ missing topology, where one can hardly distinguish between the production of a $\pi^{+}\pi^{-}$ pair off the proton and a $\pi^{0}\pi^{-}$ pair off the neutron, if only the proton and the $\pi^{-}$ in the final state were registered. Furthermore, the $\pi^{+}$ missing topology also has a strong admixture of events from the reaction $en(p)\rightarrow e^{\prime}p^{\prime}(p^{\prime})\pi^{-}$. These circumstances prevented the use of the proton missing and the $\pi^{+}$ missing topologies in this analysis.
@@ III.5.2 Fermi smearing and FSIs
Exclusive reactions off bound protons have the following features that are not present in reactions off free protons: (a) Fermi motion of the initial proton and (b) final state interactions (FSIs) of the reaction final hadrons with the spectator neutron. These features introduce some complications into the exclusive event selection, as discussed below.
@@
Since the momentum of the initial proton was not experimentally measured, this analysis uses the so-called “target-at-rest assumption” for calculation of some kinematic quantities (such as missing mass, reaction invariant mass $W$, etc.). This leads to the Fermi smearing of the corresponding experimental distributions Skorodumina et al. (2015b). To reliably identify the exclusive channel and correctly estimate the detector efficiency, a good match between the distributions of experimental events and reconstructed Monte Carlo events should be observed. This demands the simulated distributions reproduce the Fermi smearing of the experimental distributions, which implies that the effects of initial proton motion are properly included in the Monte Carlo simulation.
@@
For this reason, the Monte Carlo simulation in this analysis was performed using the TWOPEG-D Skorodumina et al. (2017a) event generator, which simulates the quasi-free process of double-pion electroproduction off a moving proton. This is an extension of TWOPEG, which is the event generator for double-pion electroproduction off the free proton Skorodumina et al. (2017b). For the TWOPEG-D version of the event generator, the Fermi motion of the initial proton is generated according to the Bonn potential Machleidt et al. (1987) and then is merged, in a natural way, into the specific kinematics of double-pion electroproduction.
@@
FSIs of the reaction final hadrons with the spectator nucleon introduce the second intrinsic feature of exclusive reactions off bound nucleons. Such interactions alter the total four-momentum of the reaction final state and therefore, introduce distortions into the distributions of some kinematic quantities (such as missing masses), thus complicating the identification of a specific exclusive channel Skorodumina et al. (2021) (see also Sec. VII.2).
@@
In contrast to the effects of the initial proton motion, which can be simulated fairly easily, the FSI effects can hardly be taken into account in the simulation because of their complex nature. The Monte Carlo simulation is hence not able to reproduce the distortions of some experimental distributions caused by FSIs with the spectator. For this reason, a proper procedure for isolation of quasi-free events from the FSI-background had to be developed.
@@
The yield of events in FSI-disturbed kinematics turned out to strongly depend on (i) the reaction invariant mass $W$ and (ii) the hadron scattering angles. The latter issue causes FSI effects to manifest themselves differently depending on the reaction topology, since the topologies have nonidentical geometrical acceptance (see Sec. VII.3). For this reason, the channel identification was performed in each topology individually, as described in the next subsections.
@@ III.5.3 Fully exclusive topology
To isolate quasi-free double-pion events in the fully exclusive topology, the distributions of the quantities determined by Eq. (2) were used. The missing momentum $P_{X}$ and the missing mass squared $M^{2}_{X[0]}$ are defined for the reaction $ep(n)\rightarrow e^{\prime}p^{\prime}(n^{\prime})\pi^{+}\pi^{-}X$, where $X$ corresponds to the undetected part.
$P_{X}$ $=$ $~{}|\overrightarrow{P}_{e}-\overrightarrow{P}_{e^{\prime}}-\overrightarrow{P}_{p^{\prime}}-\overrightarrow{P}_{\pi^{+}}-\overrightarrow{P}_{\pi^{-}}|$ (2)
$M_{X[0]}^{2}$ $=$ $~{}[P_{e}^{\mu}+P_{p}^{\mu}-P_{e^{\prime}}^{\mu}-P_{p^{\prime}}^{\mu}-P_{\pi^{+}}^{\mu}-P_{\pi^{-}}^{\mu}]^{2}$
Here $P_{i}^{\mu}$ are the four-momenta and $\overrightarrow{P_{i}}$ the three-momenta of particle $i$. Both quantities were calculated under the target-at-rest assumption, i.e. considering $P^{\mu}_{p}=(0,0,0,m_{p})$ with the proton rest mass $m_{p}$.
The quantities $P_{X}$ and $M^{2}_{X[0]}$ are unique for the fully exclusive topology as they can be calculated only if all final hadrons were registered. Figure 6 presents the distributions of $P_{X}$ (left plot) and $M^{2}_{X[0]}$ (right plot) for experimental data (in black) and Monte Carlo simulation (in blue) in a 100-MeV-wide bin in $W$.
Figure 6: Distributions of the quantities $P_{X}$ (left) and $M^{2}_{X[0]}$ (right) defined in Eq. (2) plotted for experimental data (black) and Monte Carlo simulation (blue) in one 100-MeV-wide bin in $W$. Vertical red lines indicate the cuts applied for the selection of exclusive quasi-free events. The plotted quantities as well as the values of $W$ were calculated under the target-at-rest assumption. The distributions are normalized to their maxima.
As seen in Fig. 6, the simulated $P_{X}$ distribution matches the experimental one for $P_{X}<0.2$ GeV, while for $P_{X}>0.2$ GeV the simulation underestimates the data. The mismatch mainly originates from experimental events in which final hadrons interacted with the spectator neutron. The contribution from such events cannot be reproduced by the Monte Carlo simulation as the latter does not include FSI effects. The background channels, also not included in the Monte Carlo, contribute to the mismatch, too.
The cut on the missing momentum $P_{X}$ was applied to select exclusive events in quasi-free kinematics, and the cut value was chosen to be $P_{X}=0.2$ GeV. To further clean up the sample of selected events, the cut on the missing mass squared $M^{2}_{X[0]}$ was applied complementing the cut on the missing momentum. The cuts are shown in Fig. 6 by the vertical red lines.

### v4

@@ III.2 Hadron identification
In Eq. (1), the quantity $\beta_{\text{n}}$ (which is termed the nominal value) is calculated using the particle momentum ($p_{h}$) known from the DC and the exact particle mass assumption ($m_{h}$).
To develop hadron identification cuts, experimental distributions of $\beta_{h}$ were examined as a function of the particle momentum for each TOF scintillator for each CLAS sector. This examination revealed that for some scintillation counters, the distributions were either shifted from their nominal positions or showed a double-band structure. To correct the timing information for such counters, a special procedure, described in Refs. Skorodumina et al. (2020); Skorodumina , was applied. In addition, a few counters were found to give unreliable signals; they were removed from consideration for both the experimental data and the simulation.
Figure 4: $\beta_{h}$ versus momentum distribution for positive pion candidates. The thin black curve in the middle of the event band corresponds to the nominal $\beta_{\text{n}}$ given by Eq. (1). The red curves show the applied identification cuts. Events between the red curves were treated as good pion candidates.
Figure 4 shows the $\beta_{h}$ versus momentum distribution for positive pion candidates plotted for all sectors and all reliable scintillators. The event band of the pion candidates is clearly seen. The thin black curve in the middle of the event band corresponds to the nominal $\beta_{\text{n}}$ given by Eq. (1). The red curves show the applied identification cuts. Events between the red curves were selected for further analysis. Analogous identification cuts were performed for the proton and negative pion candidates.
@@ III.3 Momentum corrections
The simulation of the CLAS detector correctly propagates particles through the media, and therefore the effect of the energy loss is already included in the estimated detector efficiency and does not impact the extracted cross sections. Nevertheless, in this study, the proton momentum magnitude was corrected for the energy loss. The simulation of the CLAS detector was used to establish the correction function, which was then applied to both experimental events and reconstructed Monte Carlo events Skorodumina et al. (2020); Skorodumina .
@@
Due to undefined origin of the above effect, it cannot be simulated, and therefore, it has become conventional for CLAS data analyses to apply a special momentum correction to the experimental data. This particular study uses the electron momentum corrections that have previously been developed and tested in the analysis of the free proton part of the “e1e” dataset at the same beam energy Fedotov et al. (2017, 2018). To establish them, the approach from Ref. Park et al. (2003), which was based on elastic kinematics, was used. The applied corrections include an electron momentum magnitude correction as well as an electron polar angle correction, which were developed for each CLAS sector individually.
@@ III.4.1 Fiducial cuts
The active detection solid angle of the CLAS detector was smaller than $4\pi$, in part due to the space occupied by the torus field coils. This is to say that the angles covered by the coils were not equipped with any detection system and therefore formed a “dead” area for particle detection Mecking et al. (2003). Additionally, the detection area was further limited in the polar angle from 8 up to 45 for electrons and up to 140 for other charged particles Mecking et al. (2003).
As was shown in different data analyses, the edges of the active detection area also do not provide a safe region for particle reconstruction, being affected by rescattering from the coils, field distortions, and similar effects. Therefore, it has become common practice to exclude these regions from consideration by applying specific fiducial cuts. This method guarantees that events accepted in the analysis include only particles detected in “safe” areas of the detector, where the acceptance is well understood.
In this study, fiducial cuts were applied for all four final-state particles ($e^{\prime}$, $p^{\prime}$, $\pi^{+}$, and $\pi^{-}$) for both experimental events and reconstructed Monte Carlo events. The analytical shapes of the cuts are similar to those used in the analysis of the free proton part of the “e1e” dataset at the same beam energy Fedotov et al. (2017, 2018), and more details can be found in Refs. Skorodumina et al. (2020); Skorodumina .
@@ III.4.2 Data quality checks
To establish data quality check cuts, the DAQ live time, as well as the yields of inclusive and elastic events normalized to the FC charge, were examined as a function of the block number. Those blocks for which these quantities demonstrated large fluctuations were excluded from the analysis. In this study, the quality check cuts are similar to those used in Refs. Fedotov et al. (2017, 2018). For more details, see also Refs. Skorodumina et al. (2020); Skorodumina .
@@ III.4.3 Vertex cut
Figure 5 presents the distributions of the electron $z$ coordinate at the interaction vertex for events from full and empty target runs (black and magenta histograms, respectively). Both distributions are normalized to the corresponding charge accumulated in the Faraday cup. The value of the vertex coordinate $z$ is corrected for the effects of the beam offset [^2] The beam offset is the deviation of the beam position from the CLAS central line $(x,y)=(0,0)$ that can lead to inaccurate determination of the vertex position. Skorodumina et al. (2020); Skorodumina . Both distributions in Fig. 5 demonstrate a well-separated peak around $z_{e^{\prime}}=2.6$ cm originating from the downstream aluminum foil. The distribution of events from the empty target runs shows two other similar peaks that correspond to the entrance and exit windows of the target cell (see also Sec. II.2).
Figure 5: Distributions of the electron $z$ coordinate at the vertex for full (black histogram) and empty (magenta histogram) target runs for CLAS sector 4. Vertical red lines show the applied cuts. Both the full and empty target distributions are normalized to the corresponding FC charge.
Empty target events were passed through the same selection procedure that was established for the liquid-deuterium data and eventually were subtracted from the latter as described in Sec. IV.5. In addition to the empty target event subtraction, a cut on the electron $z$ coordinate was applied. This cut is shown by the two vertical lines in Fig. 5: events outside these lines were excluded from the analysis.
@@ III.5.1 Reaction topologies
To identify a certain exclusive reaction, one needs to register the scattered electron and either all final hadrons or all except one. In the latter case, the four-momentum of the unregistered hadron can be deduced using energy-momentum conservation. Thus for the reaction $ep(n)\rightarrow e^{\prime}p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$ one can, in general, distinguish between four “topologies” depending on the specific combination of registered final hadrons. In this particular analysis, the following two topologies were analyzed ($X$ denotes the undetected part),
@@
  • the fully exclusive topology (all final particles are registered) $ep(n)\rightarrow e^{\prime}p^{\prime}(n^{\prime})\pi^{+}\pi^{-}X$ and
  • the $\pi^{-}$ missing topology $ep(n)\rightarrow e^{\prime}p^{\prime}(n^{\prime})\pi^{+}X$.
@@
The statistics of the fully exclusive topology are very limited, mainly because CLAS did not cover the polar angle range $0\,^{\circ}\mathrm{}<\theta_{\textrm{lab}}<8\,^{\circ}\mathrm{}$ Mecking et al. (2003). In this experiment, the presence of this forward acceptance hole mostly impacted registration of the negative particles ($e$ and $\pi^{-}$) as their trajectories were bent by the torus magnetic field towards the beam axis. This led to a constraint on the minimum achievable $Q^{2}$ for electrons and prevented registration of the majority of negative pions. As a consequence, the $\pi^{-}$ missing topology contains the dominant part of the statistics. The contribution of the fully exclusive topology to the total analyzed statistics varies from $\approx$5near the reaction threshold to $\approx$25at $W$ between 1.7 and 1.8 GeV.
@@
Besides the limited statistics, the fully exclusive topology also suffers from limited acceptance and hence from a very large number of empty cells (see Sec. V.1 for more details on the empty cells). These circumstances do not allow for any sensible cross section information to be obtained from this topology alone. The $\pi^{-}$ missing topology has a tolerable number of empty cells and large statistics and therefore serves the purpose of the cross section extraction best.
@@
For both analyzed topologies, a small admixture of the three-pion background is present. This is similar to double-pion production off free protons, where the main background channel is $ep\rightarrow e^{\prime}p^{\prime}\pi^{+}\pi^{-}\pi^{0}$. The free-proton analysis Fedotov et al. (2017, 2018) that was carried out for the same beam energy $E_{\text{b}eam}=2.039$ GeV demonstrated that although the admixture of events from this background channel becomes discernible at $W\gtrsim 1.6$ GeV, it remains negligible and well separated from double-pion events via the exclusivity cuts. For double-pion production off protons in deuterium, one more background channel can be distinguished, which is $en(p)\rightarrow e^{\prime}p^{\prime}(p^{\prime})\pi^{+}\pi^{-}\pi^{-}$, but events from this channel follow the same kinematic pattern as events from the aforementioned $ep\rightarrow e^{\prime}p^{\prime}\pi^{+}\pi^{-}\pi^{0}$ reaction.
@@
In general, two more topologies can be distinguished, i.e. the proton missing topology and the $\pi^{+}$ missing topology. Both require registration of the $\pi^{-}$ in the final state and as a result suffer from the same issues of suppressed statistics and limited acceptance as the fully exclusive topology. These two topologies are typically ignored in double-pion analyses Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017). Nevertheless, as demonstrated in a previous free-proton analysis Fedotov et al. (2017, 2018), all four reaction topologies can be used in combination, which allows for an increase in the statistics and a reduction in the number of empty cells.
@@
However, if a $\pi^{+}\pi^{-}$ pair was produced off a proton bound in deuterium, these two additional topologies turn out to be contaminated with events from other reactions. Specifically, in the proton missing topology, missing particle reconstruction fails to determine whether the pion pair was produced off the proton or off the neutron because their masses are almost identical. A similar situation occurs for the $\pi^{+}$ missing topology, where one can hardly distinguish between the production of a $\pi^{+}\pi^{-}$ pair off the proton and a $\pi^{0}\pi^{-}$ pair off the neutron, if only the proton and the $\pi^{-}$ in the final state were registered. Furthermore, the $\pi^{+}$ missing topology also has a strong admixture of events from the reaction $en(p)\rightarrow e^{\prime}p^{\prime}(p^{\prime})\pi^{-}$. These circumstances prevented the use of the proton missing and the $\pi^{+}$ missing topologies in this analysis.
@@ III.5.2 Fermi smearing and FSIs
Exclusive reactions off bound protons have the following features that are not present in reactions off free protons: (a) Fermi motion of the initial proton and (b) final-state interactions (FSIs) of the reaction final hadrons with the spectator neutron. These features introduce some complications into the exclusive event selection, as discussed below.
@@
Since the momentum of the initial proton was not experimentally measured, this analysis uses the so-called target-at-rest assumption for calculation of some kinematic quantities (such as missing masses, reaction invariant mass $W$, etc.). This leads to the Fermi smearing of the corresponding experimental distributions Skorodumina et al. (2015b, 2021) (see also Sec. IV.1). To reliably identify the exclusive channel and correctly estimate the detector efficiency, a good match between the distributions of experimental events and reconstructed Monte Carlo events should be observed. This demands the simulated distributions reproduce the Fermi smearing of the experimental distributions, which implies that the effects of the initial proton motion are properly included in the Monte Carlo simulation.
@@
For this reason, the Monte Carlo simulation in this analysis was performed using the TWOPEG-D Skorodumina et al. (2017a) event generator, which simulates a quasifree process of double-pion electroproduction off a moving proton. This is an extension of TWOPEG, which is an event generator for double-pion electroproduction off a free proton Skorodumina et al. (2017b). For the TWOPEG-D version of the event generator, the Fermi motion of the initial proton is generated according to the Bonn potential Machleidt et al. (1987) and then is naturally merged into the specific kinematics of double-pion electroproduction.
@@
FSIs of the reaction final hadrons with the spectator nucleon introduce the second intrinsic feature of exclusive reactions off bound nucleons. Such interactions alter the total four-momentum of the reaction final state and, therefore, introduce distortions into distributions of some kinematic quantities (such as missing masses), thus complicating identification of a specific exclusive channel Skorodumina et al. (2021) (see also Sec. VII.2).
@@
In contrast to the effects of the initial proton motion, which can be simulated fairly easily, FSI effects can hardly be taken into account in the simulation because of their complex nature. The Monte Carlo simulation is hence not able to reproduce the distortions of some experimental distributions caused by FSIs with the spectator. For this reason, a proper procedure for isolation of quasifree events from the FSI background had to be developed.
@@
The yield of events in FSI-disturbed kinematics turned out to strongly depend on (i) the reaction invariant mass $W$ and (ii) the hadron scattering angles. The latter issue causes FSI effects to manifest themselves differently depending on the reaction topology, since the topologies have nonidentical geometrical acceptance (see Sec. VII.3). For this reason, the channel identification was performed in each topology individually, as described in the following sections.
@@ III.5.3 Fully exclusive topology
To isolate quasifree double-pion events in the fully exclusive topology, the distributions of the quantities determined by Eq. (2) were used. The missing momentum $P_{X}$ and the missing mass squared $M^{2}_{X[0]}$ are defined for the reaction $ep(n)\rightarrow e^{\prime}p^{\prime}(n^{\prime})\pi^{+}\pi^{-}X$, where $X$ corresponds to the undetected part.
$P_{X}$ $=$ $~{}|\overrightarrow{P}_{e}-\overrightarrow{P}_{e^{\prime}}-\overrightarrow{P}_{p^{\prime}}-\overrightarrow{P}_{\pi^{+}}-\overrightarrow{P}_{\pi^{-}}|;$ (2)
$M_{X[0]}^{2}$ $=$ $~{}[P_{e}^{\mu}+P_{p}^{\mu}-P_{e^{\prime}}^{\mu}-P_{p^{\prime}}^{\mu}-P_{\pi^{+}}^{\mu}-P_{\pi^{-}}^{\mu}]^{2}.$
Here $P_{i}^{\mu}$ are the four-momenta and $\overrightarrow{P_{i}}$ the three-momenta of the particle $i$. Both quantities were calculated under the target-at-rest assumption, i.e. considering $P^{\mu}_{p}=(0,0,0,m_{p})$ with the proton rest mass $m_{p}$.
The quantities $P_{X}$ and $M^{2}_{X[0]}$ are unique for the fully exclusive topology, as they can be calculated only if all final hadrons were registered. Figure 6 presents the distributions of $P_{X}$ (left plot) and $M^{2}_{X[0]}$ (right plot) for the experimental data (in black) and the Monte Carlo simulation (in blue) in a 100-MeV-wide bin in $W$.
Figure 6: Distributions of the quantities $P_{X}$ (left) and $M^{2}_{X[0]}$ (right) defined in Eq. (2) plotted for the experimental data (black) and the Monte Carlo simulation (blue) in one 100-MeV-wide bin in $W$. Vertical red lines indicate the cuts applied to select exclusive quasifree events. The distributions are normalized to their maxima.
As seen in Fig. 6, the simulated $P_{X}$ distribution matches the experimental one for $P_{X}<0.2$ GeV, while for $P_{X}>0.2$ GeV the simulation underestimates the data. The mismatch mainly originates from experimental events in which the final hadrons interacted with the spectator neutron. The contribution from such events cannot be reproduced by the Monte Carlo simulation as the latter does not include FSI effects. The background channels, also not included in the simulation, contribute to the mismatch, too.
The cut on the missing momentum $P_{X}$ was applied to select exclusive events in quasifree kinematics, and the cut value was chosen to be $P_{X}=0.2$ GeV. To further clean up the sample of selected events, the cut on the missing mass squared $M^{2}_{X[0]}$ was applied complementing the cut on the missing momentum. The cuts are shown in Fig. 6 by the vertical red lines.